2. Compute for the Electric flux through an oval with a radius of 10m and has an electric field of 170N/C producing an angle θ=60. The electric flux is not parallel with surface and normal line.
1
Expert's answer
2020-08-04T16:00:05-0400

By, definition, for a constant electric field, the flux will be:

"\Phi = ES\cos\alpha"

where "E = 170 N\/C" is the strengh of the electric field, "S = \pi r^2" is the area of the oval, and "\alpha = 90\degree-\theta=30\degree" is an angle between the electric field lines and the normal to "S". Hence, obtain:


"\Phi = ES\cos\alpha =170\cdot \pi\cdot 10^2 \cos30\degree \approx 46252\space NC^{-1}m^2"

Answer. "46252\space NC^{-1}m^2".